OpenStudy (anonymous):

Expand the following using either the Binomial Theorem or Pascal’s Triangle. You must show your work for credit. (x - 5)^5

OpenStudy (primeralph):

Do you know how to use the factorial?

OpenStudy (anonymous):

find the fifth level of pascal's triangle for our coefficients do you know what it is?

OpenStudy (anonymous):

I've got the fifth level in front of me

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

Do I just expand it? and then simplify?

OpenStudy (anonymous):

ok fifth level is 1 5 10 10 5 1 if i recall right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so you are going to start with \[x^5\] next term will be \(5x^4(-5)^1\) or \(-25x^4\)

OpenStudy (anonymous):

next term will be \(10x^3(-5)^2\) or \(250x^3\)

OpenStudy (anonymous):

powers on \(x\) decrease each time by 1, and power of -5 increases by 1

OpenStudy (anonymous):

next term will be \(10x^2(-5)^3\) or \(-1250x^2\)

OpenStudy (anonymous):

second to last is \(5x(-5)^4=3125x\)

OpenStudy (anonymous):

and the final one is \((-5)^5=-3125\)

OpenStudy (anonymous):

put it all together and get \(x^5-25 x^4+250 x^3-1250 x^2+3125 x-3125\)

OpenStudy (anonymous):

kind of a pain, huh? lets check the answer http://www.wolframalpha.com/input/?i=%28x-5%29^5 looks good

OpenStudy (anonymous):

alright, so i put all of that as my answer?

OpenStudy (anonymous):

a HUGE pain

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes, it is all the answer all that mess

OpenStudy (anonymous):

\[(x-5)^5=x^5-25 x^4+250 x^3-1250 x^2+3125 x-3125\]

OpenStudy (anonymous):

Thank you!